- Island Remanufactured Wood Products Inc. sells its products at trade discounts of 30%, 10%. A competitor has been offering products at the same list prices but with trade discounts of 25%, 20%. Island Reman wishes to beat the competitor’s prices by offering a third trade discount. At least how big must the additional discount rate be to meet this objective
Answer #1

Let the list price be 100. For island Remanufactured the 1st discount = 30% of 100 = 30. 2nd discount = (100-30)*10% = 7. Thus selling price = 100-30-7 = 63

For the competitor 1st discount = 25% of 100 = 25. 2nd discount = (100-25)*20% = 15. Thus selling price = 100-25-15 = 60.

Now for island Remanufactured let the 3rd discount be “x”. Thus discount = 63*x%. Selling price = 63 – (63*x)%. Thus 63 – (63*x%) < 60

Or 3 < 63*x%

We get that at 4.76% discount the value comes to 3. Thus the discount value has to be at least 4.77%. If you want to round it off to the nearest % then the discount will have to be 5%.

Thus we can say that additional discount rate must at least be 5% (rounded to nearest whole %) or 4.77% (rounded to 2 decimal place)